if a brick weighs 3 pounds plus 1/2 a brick, what's the weight of 1 1/2 bricks?
Answer by richwmiller(17219)

x=1/2x+3
is another way of saying that half a brick weighs three pounds.
subtract 1/2 x from both sides
1/2x=3
there are 3 half bricks in 1 1/2 bricks
3/2=1 1/2
so 3*3=9